# A GPU-accelerated Large-scale Simulator for Transportation System Optimization Benchmarking

### v1

@@ Abstract
With the development of artificial intelligence techniques, transportation system optimization is evolving from traditional methods relying on expert experience to simulation and learning-based decision optimization methods. Learning-based optimization methods require extensive interaction with highly realistic microscopic traffic simulators for optimization. However, existing microscopic traffic simulators are computationally inefficient in large-scale scenarios and therefore significantly reduce the efficiency of the data sampling process of optimization algorithms. In addition, the optimization scenarios supported by existing simulators are limited, mainly focusing on the traffic signal control. To address these challenges and limitations, we propose the first open-source GPU-accelerated large-scale microscopic simulator for transportation system simulation. The simulator is able to iterate at 84.09Hz, which achieves 88.92 times computational acceleration in the large-scale scenario with more than a million vehicles compared to the best baseline. Based on the simulator, we implement a set of microscopic and macroscopic controllable objects and metrics to support most typical transportation system optimization scenarios. These controllable objects and metrics are all provided by Python API for ease of use. We choose five important and representative transportation system optimization scenarios and benchmark classical rule-based algorithms, reinforcement learning, and black-box optimization in four cities. The codes are available at https://github.com/tsinghua-fib-lab/moss-benchmark with the MIT License.
@@ 1 Introduction
With the increasing level of urbanization and residents’ travel demand, the urban transportation system faces heavier traffic pressure, which brings higher commuting costs, environmental pollution and other society problems, affecting the sustainable development of the city [13, 33, 30]. To alleviate the above problems, governments usually build more transportation infrastructure and optimize the existing transportation infrastructure to enhance the systems’ capacity. For instance, these transportation system optimization methods include traffic signal control, congestion pricing, etc. However, the traditional transportation system optimization process is highly dependent on the experience of experts, which is labor-heavy and often sub-optimal [22]. With the development of reinforcement learning [23, 27] and black-box optimization [10, 5], the above optimization methods have great potential for improving the transportation system. But since all of these optimization methods use extensive interaction with the environment for feedback to perform optimization, it requires that the environment be able to model the transportation system as realistically as possible and can provide feedback fast. In the field of transportation system, simulators that simulate individual motion to provide a realistic result are referred to as microscopic simulators.
@@
At present, there are several available microscopic simulators that can evaluate the efficiency of the transportation system, including SUMO [1], CityFlow [39], and CBLab [19]. However, these simulators face the following two key challenges:
@@
  • Computational inefficiency in large-scale scenarios. Since the urban transportation system is a complex system with strong direct spatial and temporal correlation between different regions, the traffic improvement in a one area may lead to congestion in another area. Therefore, the effect of transportation system optimization should be evaluated in a global city-level perspective, which poses a requirement for large-scale microscopic simulation. However, existing simulators typically use CPUs for computation. The most popular open source simulator, SUMO, is still using a single-threaded computing architecture, which significantly reduces the efficiency of the data sampling process of optimization algorithms such as reinforcement learning. Even though CityFlow and CBLab use multi-threading techniques, it still takes more than 100 seconds to simulate 1 hour in a scenario of about 100,000 vehicles. Due to the large number of environment interactions required by optimization methods especially learning-based methods, we need simulators that can simulate in large-scale scenarios with a frequency of at least 10Hz for adopting these methods into transportation system optimization.
@@
  • Limited supported optimization scenarios. In order to improve the efficiency of the transportation system, traffic management authorities usually apply a variety of transportation system optimization methods, including traffic signal control optimization, intersection lane turn assignment, tidal lane, congestion pricing, etc. If these methods can be used jointly, transportation system efficiency improvements will be further enhanced. However, existing simulators and related optimization studies usually focus on only a few of these scenarios, such as the traffic signal control optimization problem [35, 20, 37], ignoring other scenarios. This situation prevents traffic managers from fully evaluating and comparing the effectiveness of various transportation system optimization methods from microscopic like traffic signal control [42, 35, 38, 40] to macroscopic like congestion pricing [2, 4, 25]. To improve it, the simulator should be implemented to be able to support most common transportation system optimization methods and scenarios.
@@
To address the above challenges, considering the characteristics of individual independent computation in microscopic simulation matches the GPU architecture and the massive computational power of GPUs compared to CPUs, we propose the first open-source GPU-accelerated large-scale microscopic simulator for transportation system simulation. This simulator adopts a parallel-friendly design of computational flow and data partitioning, and designs an efficient indexes for sensing between vehicles. Based on these, we implement microscopic traffic simulation on CUDA and substantially improves the scale and efficiency of simulation. In the largest scenario with more than a million vehicles, this simulator is able to iterate at 84.09Hz, which is 88.92 times better than the optimal baseline. To support the optimization of various scenarios, the simulator also implements a set of microscopic and macroscopic controllable objects and metrics, and provides a Python application programming interface (API) by pybind11  [^1] https://github.com/pybind/pybind11. By combining controllable objects and metrics, we implement five typical transportation system optimization scenarios including traffic signal control, dynamic lane assignment within junctions, tidal lane control, congestion pricing, and road planning for benchmarking and evaluate the performance of classical rule-based algorithms, reinforcement learning algorithms and black-box optimization algorithms for these scenarios in 4 large cities including Beijing, Shanghai, Paris, and New York.
@@
Table 1: Comparison of microscopic simulators for transportation system. The Scale field indicates the approximate number of vehicles that can be computed by this simulator at a simulation computation frequency of 10Hz.
Simulator SUMO [1] CityFlow [39] CBLab [19] Ours
Scale (10Hz) <10000 ~130,000 ~150,000 >10,000,000
@@
In short, our contribution are two-fold. First, we propose an high-performance large-scale microscopic simulator for transportation system simulation on GPU and implement microscopic and macroscopic controllable objects and metrics to support transportation system optimization. Second, we choose and implement five typical transportation system optimization scenarios and benchmark common optimization algorithms in four cities to show the usability of our proposed simulator.
@@ 2.1 Existing Simulators for Transportation System
Existing simulators for transportation system can be divided into three categories based on the level of simplification of the simulation models: microscopic simulators, mesoscopic simulators, and macroscopic simulators. Macroscopic simulators [21, 9] typically do not consider modeling individual vehicles, but rather treat the vehicles as a fluid for using velocity and density to describe them. Mesoscopic simulators like often speed up the simulation by simplifying the vehicle motion models. For instance, MATSIM [32] use a uniform motion model with intersection waiting queues [8] to model vehicles and do not consider acceleration and deceleration. Since macroscopic and mesoscopic simulators oversimplify vehicle motion, they are not usually used for AI algorithm based transportation system optimization. Among the microscopic simulators, SUMO [1], CityFlow [39], and CBLab [19] are popular simulators for transportation system optimization. SUMO offers a rich set of controllable objects and metrics. However, due to its software architecture, SUMO can almost exclusively use one CPU core for computation, which leads to small simulation scales shown in Table 1. For CityFlow and CBLab, they both use a multi-threaded architecture for computational acceleration, which improve computational speed by about 20~30 times on 64-threaded CPUs relative to SUMO. But with city-scale simulations of at least 100,000 vehicles, it still takes minutes for them to simulate an hour, which constrains the speed of reinforcement learning algorithms to learn by interacting with the environment. Besides, in terms of controllable objects, CityFlow only provides interfaces for setting traffic signal phases and vehicle routes while CBLab adds the setting of road speed limits as an additional feature. In terms of metrics, both CityFlow and CBLab provides lane queue length and average traveling time (ATT) directly. Most of these controllable objects and metrics are designed for traffic signal optimization, and other optimization scenarios cannot be directly implemented accordingly. Overall, there is a lack of simulators that can effectively simulate and provide rich controllable objects and metrics to support transportation system optimization problems in large scale scenarios.
@@ 3 The Simulator
Figure 1: The framework and pipeline of the proposed simulator. (best view in color)
In the section, we give the overview of the design of our proposed simulator for efficient microscopic traffic simulation in large-scale urban scenarios and its interface for users as shown in Figure 1.
@@ 3.1 System Design
However, before we can simply write vehicle simulation models as CUDA code, we need to address the two problems posed by the need for vehicles to sense each other. First, in an iteration, the vehicle needs to read the position, speed, and other attributes of other vehicles as inputs to the simulation model for computing appropriate driving behaviors such as accelerating, decelerating, and changing lanes. Thereafter, the vehicle will also modify its own attributes such as position, speed, etc. based on the driving behavior of the decision. This leads to the problem of read/write conflict of vehicle data, which will affect the correctness of the simulation results. Second, the sensing behavior of a vehicle is spatially localized. Specifically, the range that the vehicle needs to sense includes only the front vehicle in the current lane and the front and rear vehicles in the adjacent lanes. Thus, implementing SIMD-friendly vehicle sensing indexes for the above retrieval task is the key to fully utilize the massive arithmetic power of the modern computational acceleration cards. Our proposed simulator designs a two-phase parallel process for read/write separation and a link-list based vehicle sensing indexes to solve the above two problems respectively. The details are described below.
@@
Two-phase Parallel Process for Read/Write Separation. In order to ensure that vehicles always correctly read the previous step’s attributes of other vehicles in the computation and to avoid interfering with the vehicle’s computation and attribute updating process, we divide the vehicle’s attributes into two partitions: snapshot and runtime. The snapshot is a read-only data partition that always saves the public attributes of the previous step for other vehicles to access. The runtime is a private and read-write partition, the attributes of which are changed after the vehicle completes its simulation calculations. In order to implement the data replication from the runtime partition to the snapshot partition, we also divide each iteration into two sequential phases, the prepare phase and the update phase. The prepare phase is used to perform vehicle data replication in parallel and update the vehicle sensing indices based on the new snapshot data. In the update phase, the vehicle performs sensing to obtain the attributes of the snapshot partitions of other vehicles and performs the car-following model [29] and the lane-changing model [12, 7] calculations to update its own runtime partition attributes. The above process effectively avoids the read/write conflict of vehicle data, which on one hand ensures the correctness of the calculation results, and on the other hand makes the calculation flow more suitable for the SIMD calculation model due to the mutex-free structure and the highly homogeneous calculation procedures.
@@
Linked-list based Vehicle Sensing Indexes. Faced with the need for SIMD-friendly spatial relative position indexing, it is not appropriate to intuitively employ a binary tree search. This is due to the fact that the process of tree searching leads to the control flow divergence problem, which significantly reduces the efficiency of the operation under the SIMD computation model. Therefore, we choose a bidirectional ordered linked list data structure to build the index. One linked list records all vehicles in order of spatial location in one lane. Each node on the linked list additionally contains two sets of pointers to the front and rear vehicles in both the left lane and the right lane, respectively. With such an index structure, vehicle sensing always requires only one pointer operation, avoiding the control flow divergence problem. Since there are usually only a small number of vehicles entering or leaving the lane at each step, and the order of the original vehicles on the lane is basically unchanged, the number of operations such as adding new nodes, deleting nodes, and reordering of the linked list during the index update process is relatively small so that the impact on the computational performance is acceptable. With this design, we address the second problem by providing a SIMD-friendly vehicle sensing indexes with low update cost for vehicle simulation model computation.
@@
To help make the simulator user-friendly, we also provide a toolchain for building simulator inputs and the simulator’s Python API.
Simulator Inputs and Toolchain. Following microscopic traffic simulation setup, the simulator inputs are map data and travel demand. The map data describes the geospatial attributes and topological relationships of road networks and the candidate traffic signal phases of junctions. Travel demand describes the vehicle’s origin, destination, departure time, and chosen route. These inputs are stored in a binary format defined by Protobuf [^2] https://protobuf.dev/. In order to facilitate the construction of simulator inputs, we have developed a toolchain available at https://github.com/tsinghua-fib-lab/mosstool. The toolchain mainly provide map building based on OpenStreetMap [^3] https://openstreetmap.org/ and real travel demand generation based on globally available public data represented by satellite imagery. By using this toolchain, users can quickly build maps, generate travel demands, and subsequently begin simulation and optimization.
@@
Python API. The simulator exposes the C interfaces as Python API via pybind11. The Python API consists of a series of initialization functions, getter functions, setter functions, and the next_step function that control the progress of the simulation. The setter functions usually provide batch versions additionally with the _batch suffix to minimize the data transfer overhead for large numbers of calls. This Python API dose not directly provide the gymnasium-style reinforcement learning environments, but rather requires users to build the environment by combining the above functions according to the need of scenarios.

### v2

@@ Abstract
With the development of artificial intelligence techniques, transportation system optimization is evolving from traditional methods relying on expert experience to simulation and learning-based decision and optimization methods. Learning-based optimization methods require extensive interactions with highly realistic microscopic traffic simulators. However, existing microscopic traffic simulators are inefficient in large-scale scenarios and thus fail to support the adoption of these methods in large-scale transportation system optimization scenarios. In addition, the optimization scenarios supported by existing simulators are limited, mainly focusing on the traffic signal control. To address these challenges, we propose the first open-source GPU-accelerated large-scale microscopic simulator for transportation system simulation and optimization. The simulator can iterate at 84.09Hz, which achieves 88.92 times computational acceleration in the large-scale scenario with 2,464,950 vehicles compared to the best baseline CityFlow. Besides, it achieves a more realistic average road speeds simulated on real datasets by adopting the IDM model as the car-following model and the randomized MOBIL model as the lane-changing model. Based on it, we implement a set of microscopic and macroscopic controllable objects and metrics provided by Python API to support typical transportation system optimization scenarios including traffic signal control, dynamic lane assignment within junctions, tidal lane control, congestion pricing, road planning, e.t.c. We choose five representative transportation system optimization scenarios and benchmark classical rule-based algorithms, reinforcement learning algorithms, and black-box optimization algorithms in four cities. These experiments effectively demonstrate the usability of the simulator for large-scale traffic system optimization. The code of the simulator is available at https://github.com/tsinghua-fib-lab/moss and the others are shown at Appendix A. In addition, we build an open-registration web platform available at https://moss.fiblab.net to support no-code trials.
@@ 1 Introduction
With the increasing level of urbanization and residents’ travel demand, the urban transportation system faces heavier traffic pressure, which brings higher commuting costs, environmental pollution and other society problems, affecting the sustainable development of the city (Kirago et al., 2022; Wang et al., 2021; Treiber et al., 2008). To alleviate the above problems, governments usually build more transportation infrastructure and optimize the existing transportation infrastructure to enhance the systems’ capacity. For instance, these transportation system optimization methods include traffic signal control, congestion pricing, etc. However, the traditional transportation system optimization process is highly dependent on the experience of experts, which is labor-heavy and often sub-optimal (McNally, 2007). With the development of reinforcement learning (Mnih et al., 2013; Schulman et al., 2017) and black-box optimization (Hansen et al., 2010; Costa & Nannicini, 2018), the above optimization methods have great potential for improving the transportation system. But since all of these optimization methods use extensive interactions with the environment for feedback to perform optimization, it requires that the environment be able to model the transportation system as realistically as possible and can provide feedback fast. In the field of transportation system, simulators that use kinetic formulations with individual motion models to calculate individual motions to provide realistic results are referred to as microscopic simulators.
@@
At present, there are several available microscopic simulators that can evaluate the efficiency of the transportation system, including SUMO (Behrisch et al., 2011), CityFlow (Zhang et al., 2019), and CBLab (Liang et al., 2023). However, these simulators face the following two key challenges:
@@
  • Computational inefficiency in large-scale scenarios. Since the urban transportation system is a complex system with strong direct spatial and temporal correlation between different regions, the traffic improvement in a one area may lead to congestion in another area. Therefore, the effect of transportation system optimization should be evaluated in a global city-level perspective, which poses a requirement for large-scale microscopic simulation. However, existing simulators typically use CPUs for computation, making it difficult to simulate city-level scenarios with a large number of individuals fast. For example, the most popular open-source simulator, SUMO, is still using a single-threaded computing architecture, which significantly reduces the efficiency of the data sampling process of optimization algorithms such as reinforcement learning. Even though CityFlow and CBLab use multi-threading techniques, it still takes more than 100 seconds to simulate 1 hour in a scenario of about 100,000 vehicles. Due to the large number of environmental interactions required by optimization methods, especially learning-based ones, existing simulators fail to support the adoption of these methods in large-scale transportation system optimization scenarios. Therefore, we need a simulator capable of simulating large-scale scenarios with about 1 million vehicles at a frequency of at least 10 Hz for supporting it.
@@
  • Limited supported optimization scenarios. In order to improve the efficiency of the transportation system, traffic management authorities usually apply a variety of transportation system optimization methods, including traffic signal control optimization, intersection lane turn assignment, tidal lane, congestion pricing, etc. If these methods can be used jointly, transportation system efficiency improvements will be further enhanced. However, existing simulators and related optimization studies usually focus on only a few of these scenarios, such as the traffic signal control optimization problem (Wei et al., 2019; Wu et al., 2023b), ignoring other scenarios. This situation prevents traffic managers from fully evaluating and comparing the effectiveness of various transportation system optimization methods from microscopic like traffic signal control (Zheng et al., 2019; Wei et al., 2019; Wu et al., 2021; Zhang et al., 2022) to macroscopic like congestion pricing (Buini et al., 2018; Pandey & Boyles, 2018). To improve it, the simulator should be implemented to be able to support more common transportation system optimization methods and scenarios.
@@
To address the above challenges, considering the characteristics of individual independent computation in microscopic simulation matches the GPU architecture and the massive computational power of GPUs compared to CPUs, we propose the first open-source GPU-accelerated large-scale microscopic simulator. This simulator adopts a parallel-friendly design of computational flow and data partitioning, and designs an efficient indexes for sensing between vehicles. Based on these, we implement microscopic traffic simulation on CUDA and substantially improves the scale and efficiency of simulation. In the largest scenario with 2,464,950 vehicles, this simulator can iterate at 84.09Hz, which is 88.92 times better than the optimal baseline CityFlow. it achieves a more realistic average road speeds than CityFlow simulated on real datasets by adopting the IDM model as the car-following model and the randomized MOBIL model as the lane-changing model. To support the optimization of various scenarios, it also implements a set of microscopic and macroscopic controllable objects and metrics, and provides a Python application programming interface (API). By combining controllable objects and metrics, we implement five typical transportation system optimization scenarios including traffic signal control, dynamic lane assignment within junctions, tidal lane control, congestion pricing, and road planning for benchmarking and evaluate the performance of classical rule-based algorithms, reinforcement learning algorithms and black-box optimization algorithms for these scenarios in 4 large cities including Beijing, Shanghai, Paris, and New York. The experiments show the usability of the simulator for large-scale traffic system optimization.
In short, our contribution are two-fold. First, we propose a realistic high-performance large-scale microscopic simulator for transportation system simulation on GPU and implement microscopic and macroscopic controllable objects and metrics to support transportation system optimization. Second, we choose and implement five typical transportation system optimization scenarios and benchmark common optimization algorithms in four cities to show the usability of our proposed simulator.
@@
Table 1: Comparison of microscopic simulators for transportation system. The Scale field indicates the approximate number of vehicles that can be computed by the simulator at a simulation computation frequency of 10Hz on an Intel(R) Xeon(R) Platinum 8462Y CPU (64 threads) and an NVIDIA GeForce RTX 4090 GPU. (Citations are left in the text to avoid out-of-width.)
Simulator SUMO CityFlow CBLab Ours
Scale (10Hz) $<10,000$ $\sim 130,000$ $\sim 150,000$ $>10,000,000$
Simulation Models Car-following Selectable [^1] Shown in the Car-Following Models and Lane-Changing Models sections at https://sumo.dlr.de/docs/Definition_of_Vehicles%2C_Vehicle_Types%2C_and_Routes.html Krauß Krauß IDM
Lane-changing Earliest Private Randomized MOBIL
@@ 2.1 Existing Simulators for Transportation System
Existing simulators for transportation system can be divided into three categories based on the level of simplification of the simulation models: microscopic simulators, mesoscopic simulators, and macroscopic simulators. Macroscopic simulators (Mahmassani, 1992; Group, ) typically do not consider modeling individual vehicles, but rather treat the vehicles as a fluid for using velocity and density to describe them. Mesoscopic simulators often speed up the simulation by simplifying the vehicle motion models. For instance, MATSIM (W Axhausen et al., 2016) use a uniform motion model with intersection waiting queues (Gawron, 1998) to model vehicles and do not consider acceleration and deceleration. Since macroscopic and mesoscopic simulators oversimplify vehicle motion, they are not usually used for AI algorithm based transportation system optimization. Among the microscopic simulators, SUMO (Behrisch et al., 2011), CityFlow (Zhang et al., 2019), and CBLab (Liang et al., 2023) shown in Table 1 are popular simulators for transportation system optimization. They all use a car-following model and a lane-change model to simulate the vehicle’s behaviors and calculate the acceleration, velocity, and position of vehicles. SUMO provides multiple simulation models for user selection and offer a rich set of controllable objects and metrics. However, due to its software architecture, SUMO can almost exclusively use one CPU core for computation, which leads to small simulation scales. For CityFlow and CBLab, they both use a multi-threaded architecture for computational acceleration, which improve computational speed by about $20\sim 30$ times on 64-threaded CPUs relative to SUMO. But with city-scale simulations of at least 100,000 vehicles, it still takes minutes for them to simulate an hour, which constrains the speed of reinforcement learning algorithms to learn by interacting with the environment. Both CityFlow and CBLab use Krauß et al. (1997) as the car-following model. For the lane-changing model, CityFlow uses an as-early-as-possible lane changing strategy, while CBLab implements its own private adaptive lane changing algorithm. Besides, in terms of controllable objects, CityFlow only provides interfaces for setting traffic signal phases and vehicle routes while CBLab adds the setting of road speed limits as an additional feature. In terms of metrics, both CityFlow and CBLab provides lane queue length and average traveling time (ATT) directly. Most of these controllable objects and metrics are designed for traffic signal optimization, and other optimization scenarios cannot be directly implemented accordingly. Overall, there is a lack of microscopic simulators that can effectively simulate and provide rich controllable objects and metrics to support transportation system optimization problems in large scale scenarios.
@@ 3 The Simulator
((a))
((b))
Figure 1: (a) The framework of the proposed simulator. (b) The linked-list of the center line for vehicle sensing by one pointer operation. (best view in color)
To address two key challenges in the simulation and optimization of large-scale urban transportation systems, we first introduce to the design of our proposed simulator for efficient microscopic traffic simulation and then discuss the microscopic and macroscopic controllable objects and metrics with their APIs to support optimization in the section. The framework of the proposed simulator is shown in Figure 1(a).
@@ 3.1 System Design
However, before we can implement vehicle simulation models as CUDA code, we need to address the two problems posed by the need for vehicles to sense each other. First, in an iteration, the vehicle needs to read the position, velocity, and other attributes of other vehicles as inputs to the simulation models for computing appropriate driving behaviors such as accelerating, decelerating, and changing lanes. Thereafter, the vehicle will also modify its own attributes such as position, velocity, etc. based on the driving behavior of the decision. This leads to the problem of read/write conflict of vehicle data, which will affect the correctness of the simulation results. Second, the range that the vehicle senses includes only the front vehicle in the current lane and the front and rear vehicles in the adjacent lanes, which is spatially localized. Thus, implementing SIMD-friendly vehicle sensing indexes for the above retrieval task is the key to fully utilize the massive arithmetic power of the modern computational acceleration cards. The two key designs of our proposed simulator to solve the above two problems respectively are described below.
@@
Two-phase Parallel Process for Read/Write Separation. In order to ensure that vehicles always correctly read the previous step’s attributes of other vehicles and to avoid the read/write conflict, we divide the vehicle’s attributes into two partitions: snapshot and runtime. The snapshot is a read-only data partition that always saves the public attributes of the previous step for other vehicles to access. The runtime is a private and read-write partition, the attributes of which are changed after the vehicle completes its simulation calculations. In order to implement the data replication from the runtime partition to the snapshot partition, we also divide each iteration into two sequential phases, the prepare phase and the update phase. The prepare phase is used to perform vehicle data replication in parallel and update the vehicle sensing indices based on the new snapshot data. In the update phase, the vehicle performs sensing to obtain the attributes of the snapshot partitions of other vehicles and performs the IDM model (Treiber et al., 2000) as the car-following model and the randomized MOBIL model (Kesting et al., 2007; Feng et al., 2021) as the lane-changing model to update its own runtime partition attributes. The above process effectively avoids the read/write conflict of vehicle data, which on one hand ensures the correctness of the calculation results, and on the other hand makes the calculation flow more suitable for the SIMD calculation model due to the mutex-free structure and the highly homogeneous calculation procedures.
@@
Linked-list based Vehicle Sensing Indexes. The key to design a SIMD-friendly spatial relative position indexing is to pick the suitable data structure. Since branching will lead to the thread divergence and reduce the execution efficiency under SIMD architecture, the adjacent element query operation with constant time complexity is required. Because vehicle lane changes will result in random insertions and deletions, such actions of the data structure should be efficient. Commonly used ordered structures include ordered vectors, ordered linked lists, and binary trees. Binary trees are discarded due to their inconstant adjacent element query. The insertion and deletion of ordered vectors will result in a large number of data movement operations, resulting in on-chip memory bandwidth pressure. Therefore, ordered linked lists are the most suitable candidates as the vehicle sensing indexes. In order to solve the query with constant time complexity for the front and rear vehicles in the adjacent lanes, we add side pointer design to the bidirectional ordered linked list. Shown in Figure 1(b), one linked list records all vehicles in order of spatial location in one lane. Each vehicle node has two pointers to the front and rear vehicle in the same lane respectively, named as main pointers. In addition, we add four pointers pointing to the front and rear vehicles in the left and right lanes respectively, named as side pointers. With such an index structure, vehicle sensing always requires only one pointer operation, avoiding the thread divergence problem. Since there are usually only a small number of vehicles entering or leaving the lane at each iteration, the number of operations including adding nodes, deleting nodes, and reordering of the linked list during the index update process is relatively small so that the impact on the computational performance is acceptable. With this design, we address the second problem by providing a SIMD-friendly vehicle sensing indexes with low update cost for vehicle simulation model computation.
@@
To make it user-friendly, we provide its Python API and a toolchain for building its inputs.
@@
Python API. The simulator exposes the C interfaces as Python API. The Python API consists of a series of initialization functions, getter functions, setter functions, and the next_step function that control the progress of the simulation. The setter functions usually provide batch versions additionally with the _batch suffix to minimize the data transfer overhead for large numbers of calls. We will mention some of the key APIs in Section 3.2 and Section 3.3 and leave most of the general APIs like get_vehicle_speeds() in the public documentation. This Python API dose not directly provide the gymnasium-style reinforcement learning environments, but rather requires users to build the environment by combining the above functions according to the need of scenarios. The benchmark open-source codes for the five scenarios provides examples of how to encapsulate the gymnasium-style reinforcement learning environments on top of the Python API.
Simulator Inputs and Toolchain. Following microscopic traffic simulation setup, the simulator inputs are map data and travel demand. The map data describes the geospatial attributes and topological relationships of road networks and the candidate traffic signal phases of junctions. Travel demand describes the vehicle’s origin, destination, departure time, and chosen route. These inputs are stored in a binary format defined by Protobuf [^2] https://protobuf.dev/. In order to facilitate the construction of simulator inputs, we have developed a open-source toolchain. The toolchain mainly provide map building based on OpenStreetMap [^3] https://openstreetmap.org/ and real travel demand generation based on globally available public data represented by satellite imagery. By using the toolchain, users can easily build maps, generate travel demands, and subsequently begin simulation and optimization.
Appendix B will introduce the system implementation details, including the execution process, vehicle model implementation, etc., and further compare the system with baselines in terms of system implementation for the interested readers.